# Data-Driven Distributed Optimization via Aggregative Tracking and Deep-Learning

### v2

@@ Abstract
In this paper, we propose a novel distributed data-driven optimization scheme. In detail, we focus on the so-called aggregative framework, a scenario in which a set of agents aim to cooperatively minimize the sum of local costs, each depending on both local decision variables and an aggregation of all of them. We consider a data-driven setup where each objective function is unknown and can be sampled at a single point per iteration (thanks to, e.g., feedback from human users or sensors). We address this scenario through a distributed algorithm combining three components: (i) a learning part leveraging neural networks to learn the local costs descent direction, (ii) an optimization routine steering the estimates according to the learned direction to minimize the global cost, and (iii) a tracking mechanism locally reconstructing the unavailable global quantities. Using tools from system theory, i.e., timescale separation and averaging theory, we formally prove that in strongly convex setups, the distributed scheme linearly converges to a neighborhood of the optimum, whose radius depends on the accuracy of the neural networks. Finally, numerical simulations validate the theoretical results.
Index Terms:
Distributed Algorithms/control, Optimization, Learning, Networks of Autonomous Agents, Data-Driven Distributed Optimization.
@@ I Introduction
The mentioned surveys do not include the so-called aggregative optimization framework which is particularly suited to model tasks arising in cooperative robotics, see the recent tutorial [4]. This novel distributed optimization scenario has been introduced by the pioneering work [5] and deals with networks of agents aiming at cooperatively minimizing sum of local functions each depending on both microscopic variables (e.g., the position of a single robot) and macroscopic ones (e.g., the spatial distribution of the entire team). Online and constrained versions of the aggregative problem are considered in [6]. Other works on distributed aggregative optimization include [7] in which a distributed method based on the Franke-Wolfe update is proposed, and [8, 9] in which momentum based algorithms are proposed. Authors in [10] propose a continuous time algorithm with non-uniform gradient gains which only requires the sign of relative state information between agents’ neighbors. While in [11] a compressed communication scheme is interlaced with the solution proposed in [5]. Aggregative optimization scenarios with uncertain environments have been explored in [12] where the distributed scheme proposed in [5] is enhanced with a Recursive Least Square (RLS) method estimating the unknown cost via feedback from the users. A similar framework is addressed in [13] where instead the learning part relies on a neural network.
@@
In this paper, we introduce DEep-Learning aggregative TrAcking (DELTA), a novel distributed data-driven optimization scheme for aggregative problems. The peculiarity of DELTA is to tackle the typical convergence issues of standard $1$-point methods based on static gradient approximations by integrating a dynamic learning strategy based on neural networks that does not increase the number of cost evaluations per iteration. More in detail, DELTA operates on a single timescale and integrates three core components: (i) a learning-oriented module, (ii) an optimization-oriented routine, and (iii) a tracking-oriented mechanism. The learning-oriented component uses local neural networks to asymptotically estimate the correct costs (and their gradients) in a data-driven fashion. Specifically, each neural network employs only a single cost evaluation per iteration taken in the neighborhood of the current local estimates. The optimization-oriented component applies an approximated distributed gradient method to the aggregative problem. Its inexactness arises from two factors: the fact that the cost functions are unknown and approximated by the networks and the need for global quantities that are locally unavailable. The tracking-oriented component addresses the second challenge by locally reconstructing the global quantities, i.e., the aggregative variable and the derivative of the global cost with respect to that variable. We analyze DELTA using tools from system theory based on timescale separation and averaging theory to formally prove that, in strongly convex settings, the algorithm linearly converges in a neighborhood of the optimal solution whose radius depends on the given accuracy capabilities of the neural networks. A preliminary version of this work focusing on the time-varying settings is available in [13]. However, in that version, the addressed scenario is significantly simpler. Indeed, at each iteration, agents can (i) access to multiple samples of their cost functions and (ii) completely train their neural networks until convergence between two consecutive optimization steps. In constrast, here, at each iteration, we can access the cost functions only at a single point and the neural networks are trained only for a single step. From a technical point of view, this translates in suitably integrating also the learning dynamics analysis in the optimization and consensus one. Additionally, in [13], the local cost functions are partially known, while in this paper, they are completely unknown.
@@ II Problem Formulation
where each $\phi_{i}:\mathbb{R}^{n_{i}}\to\mathbb{R}^{d}$ is the $i$-th contribution to $\sigma$. As already mentioned above, the functions $f_{i}$ are unknown. More in detail, we focus on data-driven scenarios where each agent $i$ can only receive a single feedback per iteration about the unknown local cost $f_{i}$. We formalizes this aspect as follows.
@@
For all $i\in\{1,\ldots,N\}$, the function $f_{i}$ is unknown but agent $i$ can receive a single feedback $f_{i}(u_{x,i},u_{\sigma,i})$ per iteration where the input pair $(u_{x,i},u_{\sigma,i})\in\mathbb{R}^{n_{i}}\times\mathbb{R}^{d}$ can be arbitrarily chosen. $\blacksquare$
@@
Now, we formallly characterize the functions appearing in problem (1). To this end, we first introduce $f_{\sigma}:\mathbb{R}^{n}\to\mathbb{R}$ to denote the overall cost function of problem (1), namely
@@
The function $f_{\sigma}$ is $\mu_{f}$-strongly convex and its gradient is $L$-Lipschitz continuous, for some $\mu_{f},L>0$. Further, for all $i\in\{1,\ldots,N\}$, $\nabla_{1}f_{i}$, $\nabla_{2}f_{i}$, and $\nabla\phi_{i}$ are Lipschitz continuous with parameters $L_{x},L_{\sigma},L_{\phi}>0$, respectively. $\blacksquare$
@@ III DELTA: Distributed Algorithm Design
$\mathrm{z}^{k+1}_{i}$ $=\sum_{j\in\mathcal{N}_{i}}a_{ij}\left(\mathrm{z}^{k}_{j}+\nabla_{2}\hat{f}_{j}\left(\mathrm{x}^{k}_{j},\mathrm{w}^{k}_{j}+\phi_{j}\left(\mathrm{x}^{k}_{j}\right),\theta_{i}^{k}\right)\right)-\nabla_{2}\hat{f}_{i}\left(\mathrm{x}^{k}_{i},\mathrm{w}^{k}_{i}+\phi_{i}\left(\mathrm{x}^{k}_{i}\right),\theta_{i}^{k}\right)$
@@ III-A Learning Update
We introduce in each agent $i$ a local neural network to reconstruct the unknown functions $f_{i}$ and their gradients $\nabla_{1}f_{i}$ and $\nabla_{2}f_{i}$ from data. We denote with $\hat{f}_{i}:\mathbb{R}^{n_{i}}\times\mathbb{R}^{d}\times\mathbb{R}^{m_{i}}\to\mathbb{R}$ the estimate of $f_{i}$, where $m_{i}\in\mathbb{N}$ is the number of parameters of the neural network. In other words, for all $i\in\{1,\ldots,N\}$, and a generic parameter $\theta_{i}\in\mathbb{R}^{m_{i}}$, we consider
@@
Once each network structure is fixed (i.e., the number of layers, the number of neurons, and activation functions are chosen), the function $\hat{f}_{i}$ is analytically available. Thus, by considering differentiable functions $\hat{f}_{i}$, we also introduce their gradients
@@
The design requirement in Assumption 4 is achieved by considering differentiable activation functions with Lipschitz continuous gradient. Examples of such activation functions include: the hyperbolic tangent, the sigmoid function, and the softplus function. $\blacksquare$
@@
Assumption 6 characterizes the ability of each neural network to reconstruct the functions $f_{i}$ and their gradients through the parameter $\epsilon$, which reflects the maximum reconstruction error when the training process is complete (i.e., when $\theta_{i}=\theta_{i}^{\star}$). Such parameter depends on the neural networks architectures, their activation functions, and the unknown functions $f_{i}$. We remark that its knowledge is not required for the algorithm implementation. $\blacksquare$
Once the learning capabilities of the considered neural networks have been enforced, we are ready to formalize the learning mechanism embedded in DELTA. An immediate but computationally expensive approach would be to apply the gradient descent method to problem (4), thereby iteratively updating an estimate $\theta_{i}^{k}\in\mathbb{R}^{m_{i}}$ according to
@@
However, the strategy in (8) would require storing the entire history of cost samples and, more importantly, recomputing an increasing number $k$ of gradients in the current estimate $\theta_{i}^{k}$. To replace the computationally prohibitive method (8), we adopt a data-driven update based on a single sample per $k$, namely
@@
Indeed, with an eye to the structure formalized in (5), we underline that (9) can be implemented by only relying on a single feedback $f_{i}(\mathrm{x}^{k}_{i}+\mathrm{d}^{k}_{x,i},\mathrm{w}^{k}_{i}+\phi_{i}(\mathrm{x}^{k}_{i})+\mathrm{d}^{k}_{\sigma,i})$ (gathered from human users or local sensors) in a neighborhood of $(\mathrm{x}^{k}_{i},\mathrm{w}^{k}_{i}+\phi_{i}(\mathrm{x}^{k}_{i}))$, i.e., of the local estimate of agent $i$ about the $i$-th solution block $x^{\star}_{i}$ and $\sigma(\mathrm{x}^{k})$, respectively.
We conclude this section by introducing $\ell:\mathbb{R}^{n}\times\mathbb{R}^{Nd}\times\mathbb{R}^{m}\to\mathbb{R}$ to compactly represent the global cost addressed by the neural networks in a decoupled manner, namely
@@ III-B Optimization Update
where $\gamma>0$ is the algorithm step size. However, the update (11) would require the knowledge of the gradients of the unknown functions $f_{i}$.
To overcome this issue, we modify the update (11) with the approximations (3), namely
$\mathrm{x}^{{k+1}}_{i}$ $=\mathrm{x}^{k}_{i}-\gamma\nabla_{1}\hat{f}_{i}\left(\mathrm{x}^{k}_{i},\sigma(\mathrm{x}^{k}),\theta_{i}^{k}\right)-\gamma\frac{\nabla\phi_{i}(\mathrm{x}^{k}_{i})}{N}\sum_{j=1}^{N}\nabla_{2}\hat{f}_{j}\left(\mathrm{x}^{k}_{j},\sigma(\mathrm{x}^{k}),\theta_{i}^{k}\right).$ (12)
Additionally, the unavailable global quantities $\sigma(\mathrm{x}^{k})$ and $\sum_{j=1}^{N}\nabla_{2}\hat{f}_{j}(\mathrm{x}^{k}_{j},\sigma(\mathrm{x}^{k}),\theta_{i}^{k})/N$ are needed in (12). We thus interlace (12) with a tracking-oriented mechanisms.
@@ III-D DELTA: Convergence Properties
To shorthen the notation, consider $\theta^{k}\!\coloneq\!\operatorname{col}(\theta^{k}_{1},\dots,\theta^{k}_{N})$, $\mathrm{x}^{k}\!\coloneq\!\operatorname{col}(\mathrm{x}^{k}_{1},\dots,\mathrm{x}^{k}_{N})$, $\mathrm{w}^{k}\!\coloneq\!\operatorname{col}(\mathrm{w}^{k}_{1},\dots,\mathrm{w}^{k}_{N})$, $\mathrm{z}^{k}\!\coloneq\!\operatorname{col}(\mathrm{z}^{k}_{1},\dots,\mathrm{z}^{k}_{N})$, $\phi(\mathrm{x}^{k})\!\coloneq\!\operatorname{col}(\phi(\mathrm{x}^{k}_{1}),\dots,\phi(\mathrm{x}^{k}_{N}))$, and for all $c\in\{1,2\}$ we define
@@ IV-A DELTA as a time-varying two-time-scale system
is invariant for system (15). Thus, we now introduce a change of coordinates which allows us to isolate the invariant portion of the state and restrict the dynamics. Moreover, within this change of variables, we also introduce the error coordinates with respect to $\theta^{\star}$ and $x^{\star}$. Formally, it is defined as
@@
which implies $(\bar{\mathrm{w}}^{k},\bar{\mathrm{z}}^{k})=(\bar{\mathrm{w}}^{0},\bar{\mathrm{z}}^{0})$ for all $k\in\mathbb{N}$. Moreover, we note that the initialization $\mathrm{w}^{k}_{i}=\mathrm{z}^{k}_{i}=0$ leads to $(\bar{\mathrm{w}}^{0},\bar{\mathrm{z}}^{0})=(0,0)$ by definition and, thus, it holds $(\bar{\mathrm{w}}^{k},\bar{\mathrm{z}}^{k})=(0,0)$ for all $k\in\mathbb{N}$. Thus, we rewrite (15) by using the new coordinates (16) and neglecting the trivial dynamics of $\bar{\mathrm{w}}^{k}$ and $\bar{\mathrm{z}}^{k}$, thus obtaining an equivalent restricted system described as
@@
$\hskip 9.0pt+\!\nabla\phi(\tilde{\mathrm{x}}^{k}\!+\!x^{\star})\hat{G}_{2}(\tilde{\mathrm{x}}^{k}\!+\!x^{\star}\!,\!R\mathrm{w}_{\perp}^{k}\!+\!\phi(\tilde{\mathrm{x}}^{k}\!+\!x^{\star}),\tilde{\theta}^{k}\!+\!\theta^{\star})+\nabla\phi(\tilde{\mathrm{x}}^{k}+x^{\star})R\mathrm{z}_{\perp}^{k}\Big{]}$ (17b)
@@
where $M\!\coloneq\!R^{\top}A-I_{Nd}$ and we adopt a more compact notation by introducing $\tilde{\ell}:\mathbb{R}^{n}\times\mathbb{R}^{\left(N-1\right)d}\times\mathbb{R}^{m}\times\mathbb{N}\to\mathbb{R}$ defined as
@@
$\tilde{\mathrm{x}}^{{k+1}}\!$ $=\!\tilde{\mathrm{x}}^{k}\!-\!\gamma\!\left[\hat{G}_{1}\!\left(\tilde{\mathrm{x}}^{k}\!+\!x^{\star}\!,R\mathrm{w}_{\perp}^{k}\!+\!\phi\!\left(\tilde{\mathrm{x}}^{k}\!+\!x^{\star}\right)\!,\tilde{\theta}^{k}\!+\!\theta^{\star}\right)\right.+\!\nabla\phi\!\left(\tilde{\mathrm{x}}^{k}\!\!+\!x^{\star}\right)\!\hat{G}_{2}\!\left(\!\tilde{\mathrm{x}}^{k}\!\!+\!x^{\star}\!,R\mathrm{w}_{\perp}^{k}\!+\!\phi\!\left(\tilde{\mathrm{x}}^{k}\!\!+\!x^{\star}\right)\!,\tilde{\theta}^{k}\!+\!\theta^{\star}\!\right)$
$\hskip 9.0pt+\!\nabla\phi\!\left(\tilde{\mathrm{x}}^{k}\!+\!x^{\star}\right)\!R\mathrm{z}_{\perp}^{k}\!-\!\mathrm{p}_{1}\!\left(\tilde{\mathrm{x}}^{k}\!+\!x^{\star}\right)\!-\!\nabla\phi\!\left(\tilde{\mathrm{x}}^{k}\!+\!x^{\star}\right)\!\frac{\mathbf{1}\mathbf{1}^{\top}}{N}\mathrm{p}_{2}\!\left(\tilde{\mathrm{x}}^{k}\!+\!x^{\star}\right)\!\Big{]}\!\!$ (18b)
@@
where $\mathrm{p}_{c}(x)\!\coloneq\!\operatorname{col}(\mathrm{p}_{c,1}(x),\dots,\mathrm{p}_{c,N}(x))$ for all $c\!\in\!\{1,2\}$, while $\mathrm{p}_{\ell}\left(x,k\right)\!\coloneq\!\operatorname{col}(\mathrm{p}_{\ell,1}\!\left(x,k\right),\dots,\mathrm{p}_{\ell,N}\!\left(x,k\right))$. Essentially, differently from the real system (17), the presence of the fictitious $\mathrm{p}_{1}$, $\mathrm{p}_{2}$, and $\mathrm{p}_{\ell}$ makes $(0,0,-R^{\top}\phi\left(x^{\star}\right),-R^{\top}\hat{G}_{2}\left(x^{\star},\mathbf{1}\sigma\left(x^{\star}\right),\theta^{\star}\right))$ an equilibrium point of (18) (cf. Assumption 6). In detail, we note that the function $h:\mathbb{R}^{m}\times\mathbb{R}^{n}\to\mathbb{R}^{2(N-1)d}$ defined as
$h\left(\theta,x\right)\coloneq\begin{bmatrix}-R^{\top}\phi\left(x\right)\\ -R^{\top}\hat{G}_{2}\left(x,\mathbf{1}\sigma\left(x\right),\theta\right)\end{bmatrix}$ (19)
parametrizes the equilibrium points of the subsystems (18c)-(18d) in the states $(\theta,\mathrm{x})$ of the subsystems (18a)-(18b). Further, we note that the variations of (18a)-(18b) can be arbitrarily reduced via $\gamma$. From these observations, we interpret system (18) as a time-varying two-time-scale system, where (18a)-(18b) is the so-called slow dynamics, while (18c)-(18d) is the so-called fast part. According to this view, we now proceed by analyzing the boundary-layer and reduced system associated to (18).

### v3

@@ Abstract
In this paper, we propose a novel distributed data-driven optimization scheme. In particular, we focus on the so-called aggregative framework, namely, the scenario in which a set of agents aim to cooperatively minimize the sum of local costs, each depending on both local decision variables and an aggregation of all of them. We consider a data-driven setup in which each objective function is unknown and can be only sampled at a single point per iteration (thanks to, e.g., feedback from human users or physical sensors). We address this scenario through a distributed algorithm that combines three key components: (i) a learning part that leverages neural networks to learn the local cost functions descent direction, (ii) an optimization routine that steers the estimates according to the learned direction to minimize the global cost, and (iii) a tracking mechanism that locally reconstructs the unavailable global quantities. By using tools from system theory, i.e., timescale separation and averaging theory, we formally prove that, in strongly convex setups, the overall distributed strategy linearly converges in a neighborhood of the optimal solution whose radius depends on the given accuracy capabilities of the neural networks. Finally, we corroborate the theoretical results with numerical simulations.
@@ I Introduction
The mentioned surveys do not include the so-called aggregative optimization framework which is particularly suited to model tasks arising in cooperative robotics, see the recent tutorial [35]. This novel distributed optimization scenario has been introduced by the pioneering work [20] and deals with networks of agents aiming at cooperatively minimizing the sum of local functions each depending on both microscopic variables (e.g., the position of a single robot) and macroscopic ones (e.g., the spatial distribution of the entire team). Online and constrained versions of the aggregative problem are considered in [21]. Other works on distributed aggregative optimization include [36] in which a distributed method based on the Frank-Wolfe update is proposed, and [6, 37] in which momentum-based algorithms are proposed. Authors in [8] propose a continuous time algorithm with non-uniform gradient gains which only requires the sign of relative state information between agents’ neighbors. While in [7] a compressed communication scheme is interlaced with the solution proposed in [20]. Aggregative optimization scenarios with uncertain environments have been explored in [5] where the distributed scheme proposed in [20] is enhanced with a Recursive Least Square (RLS) method estimating the unknown cost via feedback from the users. A similar framework is addressed in [3] where instead the learning part relies on a neural network.
@@
In this paper, we introduce DEep-Learning aggregative TrAcking (DELTA), a novel distributed data-driven optimization scheme for aggregative problems. The peculiarity of DELTA is to tackle the typical convergence issues of standard $1$-point methods based on static gradient approximations by integrating a dynamic learning strategy based on neural networks that does not increase the number of cost evaluations per iteration. More in detail, DELTA operates on a single timescale and integrates three core components: (i) a learning-oriented module, (ii) an optimization-oriented routine, and (iii) a tracking-oriented mechanism. The learning-oriented component uses local neural networks to asymptotically estimate the correct costs (and their gradients) in a data-driven fashion. Specifically, each neural network employs only a single cost evaluation per iteration taken in the neighborhood of the current local estimates. The optimization-oriented component applies an approximated distributed gradient method to the aggregative problem. Its inexactness arises from two factors: the fact that the cost functions are unknown and approximated by the networks and the need for global quantities that are locally unavailable. The tracking-oriented component addresses the second challenge by locally reconstructing the global quantities, i.e., the aggregative variable and the derivative of the global cost with respect to that variable. We analyze DELTA using tools from system theory based on timescale separation and averaging theory to formally prove that, in strongly convex settings, the algorithm linearly converges in a neighborhood of the optimal solution whose radius depends on the given accuracy capabilities of the neural networks. A preliminary version of this work focusing on the time-varying settings is available in [3]. However, in that version, the addressed scenario is significantly simpler. Indeed, at each iteration, agents can (i) access multiple samples of their cost functions and (ii) completely train their neural networks until convergence between two consecutive optimization steps. In contrast, here, at each iteration, we can access the cost functions only at a single point and the neural networks are trained only for a single step. From a technical point of view, this translates in suitably integrating also the learning dynamics analysis in the optimization and consensus one. Additionally, in [3], the local cost functions are partially known, while in this paper, they are completely unknown.
@@ II Problem Formulation
where each $\phi_{i}:\mathbb{R}^{n_{i}}\to\mathbb{R}^{d}$ is the $i$-th contribution to $\sigma$. As already mentioned above, the functions $f_{i}$ are unknown. More in detail, we focus on data-driven scenarios where each agent $i$ can only receive a single feedback per iteration about the unknown local cost $f_{i}$. We formalize this aspect as follows.
@@
For all $i\in\{1,\ldots,N\}$, $f_{i}$ is unknown but agent $i$ can receive a single feedback $f_{i}(u_{x,i},u_{\sigma,i})$ per iteration where $(u_{x,i},u_{\sigma,i})\in\mathbb{R}^{n_{i}}\times\mathbb{R}^{d}$ can be arbitrarily chosen. $\blacksquare$
@@
Now, we formally characterize the functions appearing in problem (1). To this end, we first introduce $f_{\sigma}:\mathbb{R}^{n}\to\mathbb{R}$ to denote the overall cost function of problem (1), namely
@@
The function $f_{\sigma}$ is $\mu_{f}$-strongly convex and its gradient is $L$-Lipschitz continuous, for some $\mu_{f},L>0$. Further, for all $i\in\{1,\ldots,N\}$, $\nabla_{1}f_{i}$, $\nabla_{2}f_{i}$, and $\phi_{i}$ are Lipschitz continuous with parameters $L_{x},L_{\sigma},L_{\phi}>0$, respectively. $\blacksquare$
@@ III DELTA: Distributed Algorithm Design
$\mathrm{z}^{k+1}_{i}$ $=\sum_{j\in\mathcal{N}_{i}}a_{ij}\left(\mathrm{z}^{k}_{j}+\nabla_{2}\hat{f}_{j}\left(\mathrm{x}^{k}_{j},\mathrm{w}^{k}_{j}+\phi_{j}\left(\mathrm{x}^{k}_{j}\right),\theta_{j}^{k}\right)\right)-\nabla_{2}\hat{f}_{i}\left(\mathrm{x}^{k}_{i},\mathrm{w}^{k}_{i}+\phi_{i}\left(\mathrm{x}^{k}_{i}\right),\theta_{i}^{k}\right)$
@@ III-A Learning Update
We introduce in each agent $i$ a local neural network to reconstruct $f_{i}$ and the gradients $\nabla_{1}f_{i}$ and $\nabla_{2}f_{i}$ from data. Let $\hat{f}_{i}:\mathbb{R}^{n_{i}}\times\mathbb{R}^{d}\times\mathbb{R}^{m_{i}}\to\mathbb{R}$ be the estimate of $f_{i}$, where $m_{i}\in\mathbb{N}$ is the number of parameters of the neural network. In other words, for a generic parameter $\theta_{i}\in\mathbb{R}^{m_{i}}$, we consider
@@
Once each network structure is fixed (i.e., the number of layers, the number of neurons, and the activation functions are chosen), the function $\hat{f}_{i}$ is analytically available. Thus, by considering differentiable functions $\hat{f}_{i}$, we also introduce their gradients
@@
The requirement in Assumption 4 is satisfied by choosing differentiable activation functions with Lipschitz‐continuous gradients, such as tanh, sigmoid, or softplus. $\blacksquare$
@@
Assumption 6 specifies each neural network’s ability to approximate the functions $f_{i}$ and their gradients. This capability is captured by the parameter $\epsilon$, the maximum reconstruction error at convergence of the training process (i.e., when $\theta_{i}=\theta_{i}^{\star}$). Its value depends on the network design and the unknown $f_{i}$, but it is not required by the algorithm. $\blacksquare$
Once the capabilities of the neural network have been enforced, we can formalize the learning mechanism embedded in DELTA. An immediate but computationally expensive approach would be to apply the gradient descent method to problem (4), iteratively updating an estimate $\theta_{i}^{k}\in\mathbb{R}^{m_{i}}$ as
@@
However, the strategy in (8) would require storing the entire cost samples history and, more importantly, computing an increasing number $k$ of gradients in the current estimate $\theta_{i}^{k}$. To replace the computationally prohibitive method (8), we adopt a data-driven update based on a single sample per $k$, namely
@@
Indeed, with an eye to the structure formalized in (5), we underline that (9) can be implemented by only relying on a single feedback $f_{i}(\mathrm{x}^{k}_{i}+\mathrm{d}^{k}_{x,i},\mathrm{w}^{k}_{i}+\phi_{i}(\mathrm{x}^{k}_{i})+\mathrm{d}^{k}_{\sigma,i})$ in a neighborhood of $(\mathrm{x}^{k}_{i},\mathrm{w}^{k}_{i}+\phi_{i}(\mathrm{x}^{k}_{i}))$, i.e., of the local estimate of agent $i$ about the $i$-th solution block $x^{\star}_{i}$ and $\sigma(\mathrm{x}^{k})$, respectively. Finally, we compactly represent the global cost addressed by the neural networks via $\ell:\mathbb{R}^{n}\times\mathbb{R}^{Nd}\times\mathbb{R}^{m}\to\mathbb{R}$ defined as
@@ III-B Optimization Update
where $\gamma>0$ is the step size. Nevertheless, update (11) requires the gradients of the unknown $f_{i}$. We therefore modify it using the approximations (3), namely
$\mathrm{x}^{{k+1}}_{i}$ $=\mathrm{x}^{k}_{i}-\gamma\nabla_{1}\hat{f}_{i}\left(\mathrm{x}^{k}_{i},\sigma(\mathrm{x}^{k}),\theta_{i}^{k}\right)-\gamma\frac{\nabla\phi_{i}(\mathrm{x}^{k}_{i})}{N}\sum_{j=1}^{N}\nabla_{2}\hat{f}_{j}\left(\mathrm{x}^{k}_{j},\sigma(\mathrm{x}^{k}),\theta_{j}^{k}\right).$ (12)
Additionally, the unavailable global quantities $\sigma(\mathrm{x}^{k})$ and $\sum_{j=1}^{N}\nabla_{2}\hat{f}_{j}(\mathrm{x}^{k}_{j},\sigma(\mathrm{x}^{k}),\theta_{j}^{k})/N$ are needed in (12). We thus interlace (12) with a tracking-oriented mechanism.
@@ III-D DELTA: Convergence Properties
To shorten the notation, consider $\theta^{k}\!\coloneq\!\operatorname{col}(\theta^{k}_{1},\dots,\theta^{k}_{N})$, $\mathrm{x}^{k}\!\coloneq\!\operatorname{col}(\mathrm{x}^{k}_{1},\dots,\mathrm{x}^{k}_{N})$, $\mathrm{w}^{k}\!\coloneq\!\operatorname{col}(\mathrm{w}^{k}_{1},\dots,\mathrm{w}^{k}_{N})$, $\mathrm{z}^{k}\!\coloneq\!\operatorname{col}(\mathrm{z}^{k}_{1},\dots,\mathrm{z}^{k}_{N})$, $\phi(\mathrm{x}^{k})\!\coloneq\!\operatorname{col}(\phi(\mathrm{x}^{k}_{1}),\dots,\phi(\mathrm{x}^{k}_{N}))$, and for all $c\in\{1,2\}$ we define
@@ IV-A DELTA as a time-varying two-time-scale system
is invariant for system (15). Thus, we introduce a change of coordinate to isolate the invariant part of the state and define the error coordinates relative to $\theta^{\star}$ and $x^{\star}$. Formally, it is
@@
which implies $(\bar{\mathrm{w}}^{k},\bar{\mathrm{z}}^{k})=(\bar{\mathrm{w}}^{0},\bar{\mathrm{z}}^{0})$ for all $k\in\mathbb{N}$. Further, the initialization $\mathrm{w}^{k}_{i}=\mathrm{z}^{k}_{i}=0$ leads to $(\bar{\mathrm{w}}^{0},\bar{\mathrm{z}}^{0})=(0,0)$ and, thus, it holds $(\bar{\mathrm{w}}^{k},\bar{\mathrm{z}}^{k})=(0,0)$ for all $k\in\mathbb{N}$. Hence, we rewrite (15) in the new coordinates (16) and neglect the trivial dynamics of $\bar{\mathrm{w}}^{k}$ and $\bar{\mathrm{z}}^{k}$, thus obtaining an equivalent restricted system
@@
$\hskip 9.0pt+\!\nabla\phi(\tilde{\mathrm{x}}^{k}\!+\!x^{\star})\hat{G}_{2}(\tilde{\mathrm{x}}^{k}\!+\!x^{\star}\!,\!R\mathrm{w}_{\perp}^{k}\!+\!\phi(\tilde{\mathrm{x}}^{k}\!+\!x^{\star}),\tilde{\theta}^{k}\!+\!\theta^{\star})$
$\hskip 9.0pt+\nabla\phi(\tilde{\mathrm{x}}^{k}+x^{\star})R\mathrm{z}_{\perp}^{k}\Big]$ (17b)
@@
where $M\coloneq R^{\top}(A-I_{Nd})$ and we introduce the compact notation $\tilde{\ell}:\mathbb{R}^{n}\times\mathbb{R}^{(N-1)d}\times\mathbb{R}^{m}\times\mathbb{N}\to\mathbb{R}$, defined as
@@
$\tilde{\mathrm{x}}^{{k+1}}\!$ $=\!\tilde{\mathrm{x}}^{k}\!-\!\gamma\!\left[\hat{G}_{1}\!\left(\tilde{\mathrm{x}}^{k}\!+\!x^{\star}\!,R\mathrm{w}_{\perp}^{k}\!+\!\phi\!\left(\tilde{\mathrm{x}}^{k}\!+\!x^{\star}\right)\!,\tilde{\theta}^{k}\!+\!\theta^{\star}\right)\right$.
$\hskip 9.0pt+\!\nabla\phi\!\left(\tilde{\mathrm{x}}^{k}\!\!+\!x^{\star}\right)\!\hat{G}_{2}\!\left(\!\tilde{\mathrm{x}}^{k}\!\!+\!x^{\star}\!,R\mathrm{w}_{\perp}^{k}\!+\!\phi\!\left(\tilde{\mathrm{x}}^{k}\!\!+\!x^{\star}\right)\!,\tilde{\theta}^{k}\!+\!\theta^{\star}\!\right)$
$\hskip 9.0pt+\!\nabla\phi\!\left(\tilde{\mathrm{x}}^{k}\!+\!x^{\star}\right)\!R\mathrm{z}_{\perp}^{k}$
$\hskip 9.0pt\!-\!\mathrm{p}_{1}\!\left(\tilde{\mathrm{x}}^{k}\!+\!x^{\star}\right)\!-\!\nabla\phi\!\left(\tilde{\mathrm{x}}^{k}\!+\!x^{\star}\right)\!\tfrac{\mathbf{1}\mathbf{1}^{\top}}{N}\mathrm{p}_{2}\!\left(\tilde{\mathrm{x}}^{k}\!+\!x^{\star}\right)\!\Big]\!\!$ (18b)
@@
where $\mathrm{p}_{c}(x)\!\coloneq\!\operatorname{col}(\mathrm{p}_{c,1}(x),\dots,\mathrm{p}_{c,N}(x))$ for all $c\!\in\!\{1,2\}$, while $\mathrm{p}_{\ell}\left(x,k\right)\!\coloneq\!\operatorname{col}(\mathrm{p}_{\ell,1}\!\left(x,k\right),\dots,\mathrm{p}_{\ell,N}\!\left(x,k\right))$. Due to $\mathrm{p}_{1}$, $\mathrm{p}_{2}$, and $\mathrm{p}_{\ell}$, $(0,0,-R^{\top}\phi(x^{\star}),-R^{\top}\hat{G}_{2}(x^{\star},\mathbf{1}\sigma(x^{\star}),\theta^{\star}))$ is an equilibrium of (18) (cf. Assumption 6). In detail, the equilibrium points of the subsystems (18c)-(18d) are parametrized in the states $(\theta,\mathrm{x})$ of (18a)-(18b) through $h:\mathbb{R}^{m}\times\mathbb{R}^{n}\to\mathbb{R}^{2(N-1)d}$ defined as
$h\left(\theta,x\right)\coloneq\begin{bmatrix}-R^{\top}\phi\left(x\right)\\ -R^{\top}\hat{G}_{2}\left(x,\mathbf{1}\sigma\left(x\right),\theta\right)\end{bmatrix}$. (19)
Further, the variations of (18a)-(18b) can be arbitrarily reduced via $\gamma$. From these observations, we interpret system (18) as a time-varying two-time-scale system, where (18a)-(18b) is the so-called slow dynamics, while (18c)-(18d) is the so-called fast part. According to this view, we now proceed by analyzing the boundary-layer and reduced system associated to (18).